x(.10)-[.12(10000-x)]=540
Answer by Fombitz(32388)

x(.10)-[.12(10000-x)]=540
First off, let’s get rid of the decimals by multiplying both sides by 100.
100(x(.10)-[.12(10000-x)])=100(540) Use the distributive property
100x(.10)-100[.12(10000-x)]=54000
100x(.10)-100[.12(10000-x)]=54000
10x-[12(10000-x)]=54000
Start in the parentheses and work your way out. Use the distributive property.
10x-[12(10000)-12x)]=54000
10x-[120000-12x)]=54000
10x+(-1)[120000-12x)]=54000 Use the distributive property again.
10x+(-120000)+12x=54000
22x+(-120000)+120000=54000+120000 Use additive inverse of (-120000) or 120000.
22x=174000
22x/22=174000/22 Use multiplicative inverse of (22) or 1/(22).
x=174000/22=87000/11 Simplify.
x~7909.1 Approximate answer.